One employee makes 20 boxes in 60 minutes. Another employee makes 25 boxes in 60 minutes. How many boxes can they make together in 40 minutes?
Answer by LinnW(1048)

f = first employee boxes per minute
s = second employee boxes per minute
rate for first employee in boxes per minute.
20 boxes / 60 minutes = f/1
20 = 60s
divide each side by 60
20/60 = f
f = 1/3 boxes per minute
Similarly
25/60 = s
s = (5*5)/(12*5)
s = 5/12 boxes per minute
( combined rate r + s )* 40 = boxes in 40 minutes
( 1/3 + 5/12 )40
( 4/12 + 5/12 )40
( 9/12 )40
( 3/4 )40
( 120/4 )
30 boxes